OpenStudy (death47):

What is the area of the trapezoid? A.) 96 square feet B.)240 square feet C.)480 square feet D.)52 square feet I think it is B.)240ft

OpenStudy (purple_pink):

Area=1/2 h(b1+b2) - Trapezoid

OpenStudy (purple_pink):

use that formula^

rebeccaxhawaii (rebeccaxhawaii):

\[A=\frac{ a+b }{ 2 }h\]

jabez177 (jabez177):

\[A=\frac{ a+b }{ 2 }h\]

rebeccaxhawaii (rebeccaxhawaii):

h=12 a=24 b=16

OpenStudy (purple_pink):

i was 1st to give that formula :D

rebeccaxhawaii (rebeccaxhawaii):

plug it in.

OpenStudy (death47):

Alright, if I'm doing this right it should be 240: 24+16/2*12=240 ?

OpenStudy (purple_pink):

24+16=40 40x12=480 480÷2=240

OpenStudy (purple_pink):

yeah that's correct! yayyy! *claps*

rebeccaxhawaii (rebeccaxhawaii):

Correct c:

OpenStudy (death47):

^-^ thank you guys. now who do I give the medal to? ;~;

OpenStudy (purple_pink):

i'm fine w/ a fan :D you can medal becca

OpenStudy (death47):

Alright, done and le done.

OpenStudy (purple_pink):

Thank you!